# A blueprint for large-scale quantum-network deployments

### v1

@@ I Introduction
The principles of quantum mechanics imply that the measurement of a quantum signal alters it. Passive networks, without active elements such as amplifiers, must be used for its transmission. The propagation of signals at the single quantum level in a physical media—typically the optical fibre—entails an absorption probability that dramatically limits its range. If that medium is shared, other quantum or classical signals may limit the range even more. In the case of optical fibres, deployment is typically the most expensive part and co-propagation is essential. Until quantum repeaters are developed [3], it is necessary for QKD networks to handle these limitations.
Equally relevant to the success of quantum communications is the provision of the quantum-enhanced services. This involves the joint operation and management of the quantum and classical resources, which is an enabling technique of low-level ideas, like sharing the underlying physical infrastructure. Other important aspects for widespread deployment are the interoperability of deployed systems, standardisation, certification of quantum networks and devices, compliance with regulations, and quality assurance. MadQCI [4] is designed to be heterogeneous at the quantum as well as the classical network level, to comprehensively address these problems. It is also large and complex enough to significantly advance the state of the art. To date, with three years of continuous operation, MadQCI is the most complex, largest, and longest running testbed for quantum-communication technologies in Europe. To realise a diverse network and perform a comprehensive investigation, multiple technologies from various vendors were installed. The 26 QKD modules came from three different manufacturers, some using continuous-variable QKD (CV-QKD) and others discrete-variable QKD (DV-QKD). Table I summarises the main features of the network.
@@
Domain Nodes QKD links Total link length QKD modules Encryp.
@@
The complexity of the network was tackled using several administrative domains. The infrastructure was hosted in production premises of two network operators: Telefónica Innovación Digital (TID), the research and development division of the incumbent telecommunications operator in Spain, and the research and education network in the region of Madrid, REDIMadrid (RM). In addition, a third domain was used to handle the testing of a switched QKD network. This specific sub-network is referred to as the “co-located domain” and contains all the devices [5] that support switching. The network configuration resembles production networks—including the inter-domain border link—re-using already installed resources and adding new ones as required.
Regardless of the domain, the operation and management of the overall quantum network was based on a software-defined approach [4] to quantum networking. In the software-defined networking (SDN) paradigm, forwarding-plane resources are programmable and expose their capabilities through common and open interfaces [6]. In this way, the operation and management system not only configures and commands the resources, but also achieves the overall programming of the network for application-specific behaviour and performance. This design is very flexible and allows an easier integration of quantum resources in existing classical networks. In MadQCI, this paradigm enabled in a transparent way the operation and management of both the quantum and classical equipment required for quantum-secure communications and services.
As a result, cryptographic keys could be distributed between any pair of nodes in the network, independently of the QKD manufacturer and network operator. These capabilities were exposed to applications through standardised interfaces, which benefited from the programmability of the network as a whole. The outcome allowed the execution of a wide variety of use cases [4, 7, 8] fed with quantum-distributed keys. Specific hardware was installed to provide encryption at OSI levels 1 and 2. Software applications were used for test on higher OSI layers, such as an IPsec implementation [9] and one based on TLS [10].
The paper is organised as follows. In the next sections, the related work and an overview about quantum and optical communications is outlined and the coexistence and integration approach of classical and quantum optical networks is discussed. Then, a detailed description of the solutions implemented in MadQCI and its main performance are given to serve as a blueprint for large-scale quantum network deployments. The paper finishes with conclusions and an outlook.
@@ II Related work
Quantum cryptography has been the first major technological expression of quantum communication and QKD grew around the idea of using optical communications and networks since its earliest days. Examples include the Boston testbed [11], which already involved optical networking techniques such as switching or dense wavelength division multiplexing (DWDM). Other well-known examples are the Vienna [12] and Tokyo [13] quantum networks, which used dark optical fibres deployed in the field. The first iterations of MadQCI also date back to that time [14], along with others such as the Hefei testbed [15].
Most of these practical exercises focused on showing the capabilities of quantum-communication developments, rather than on how to install these into existing infrastructures [16, 17, 18, 19, 20]. In contrat, the efforts in MadQCI were aligned with this last goal since its early days. Novel optical networking approaches to quantum communication were proposed [21] and it was shown how a pioneering commitment to a software-defined approach to quantum networking enables a fast and flexible convergence of quantum and classical communication in production environments [22, 23]. Furthermore, the MadQCI team is a leading actor in the standardisation of QKD communication at ETSI ISG QKD [24], which promotes a dialogue between industrial actors such as manufacturers and network operators.
In parallel, numerous technological improvements over the last decade have brought QKD to commercial status. Deployments such as the 2,000 km Beijing-Shanghai QKD link [25] and the 4,600 km long QKD space segment [26] have demonstrated the feasibility of quantum secured broad backbone networks. Other examples demonstrate the penetration of these technologies in specific niches, such as the ability of the Hefei quantum network to efficiently encrypt its institutional telephone traffic [27] or the efforts in Bristol to securely transport subscriber databases in 5G [28].
Since 2019, a broad effort is underway to establish he European Quantum Communications Infrastructure (EuroQCI), a pan-European quantum network based on the results of several projects funded by Quantum Flagship [29] and Digital Europe Programme [30] research programmes, among others. Examples of these projects are OpenQKD [7], CIVIQ [31], and the current QSNP [32], which strongly contribute to the discussions in this paper. Currently, there are several efforts to deploy quantum communication infrastructures (QCI) in Europe. A list of these can be found in [33].
@@ III Quantum and optical communications overview
This section introduces some topics in telecommunications and quantum technologies, particularly those transmitting in optical fibres.
@@ III-A Production networks
When considering the embodiment of quantum technologies in optical networks, it is necessary to build on a suitable model, which may differ depending on the nature of the requirements [36].
The current networking landscape relies on several heterogeneous communication systems with different design, technology, and service fundamentals. Besides, each of them has a different degree of interoperability, often limited by the commercial strategy of the vendors. Examples are the more traditional Ethernet-based local area networks; the passive optical networks (PONs) for home access; or the optical networks for aggregation and backhaul, as Fig. 1 depicts. All operate over optical fibres and can carry the same type of traffic. However, they have different requirements for the deployment of new systems. In addition, network operators and service providers must comply with legal and regulatory requirements—e.g. critical infrastructure protection, lawful interception, and secrecy of communications—that prevent the deployment of novel devices without certification and industry-wide standards. The deployment of quantum communication adds additional requirements to these already overloaded designs, which have to be integrated to deliver cutting-edge services.
@@
The different industrial and commercial approaches should also be taken into account when modelling the optical networks. Network operators focus their activities on the transport of information between their subscribers. Thus, their network design must deal with the complexity of equipment spread across customer sites and central offices, while assuring the aggregation and transport of multi-service traffic. Whereas, in the case of digital service providers, the complexity lies in their multi-tenant, distributed data centres, which are interconnected “over-the-top” of the network operator, abstracting the complexity of many parts of the network. The conception of quantum communications and the service they provide can be fundamentally different in these scenarios, and involves variables ranging from regulatory issues to how or to whom to bill for these services.
A description of a network can be achieved by explaining its network functionalities, usually classified in functional planes—i.e., the forwarding, control, management, application, etc. planes—. The previously explained context bounds how, when, and where those functionalities are executed. In this regard, a production network cannot be modeled as a naive graph of connected nodes. It is common for forwarding-plane functionalities, i.e., those functions that enable the transport of information over the different transmission media, to reside at all nodes of a network. However, the majority of the control-plane functionalities may reside in a single centralised component. In this case, the unavailability of a single node may compromise the connectivity of the entire graph. Furthermore, other elements, such as the facilities that physically host the forwarding devices and provide security and privacy to the system, are also functional to the network. Therefore, a more appropriate model should include which network functionalities are performed on which element, with a focus on the specific end use or application.
@@ III-B Optical networks
Finally, several visions or generations coexist in the industry to achieve a higher or lower reconfigurability of lightpaths, as also depicted in Fig. 1. A common approach is called “FOADM”, which stands for fixed OADM. In these designs, elements such as fixed filters and multiplexers prevail, i.e., without the possibility of changing their frequency selectivity. Re-tunable transceivers may be used, depending on the connectivity requirements in the assembly. The “ROADM” approach uses precisely reconfigurable OADMs or analogous elements. This means that the optical mesh can be reconfigured by means of frequency selective elements to establish new lightpaths on demand. The degree of meshing may not be all-with-all, and hybrid approaches are possible—e.g., FOADM in the access network and ROADM in the backbone, or vice versa—.
@@ III-C Quantum key distribution and cryptographic applications
QKD provides synchronised sources of symmetric cryptographic keys between two remote systems [1, 2]. This technique relies on quantum physics to deliver information-theoretically secure (ITS) keys to a final cryptographic application [37], as Fig. 2 shows. Indeed, an ITS cryptosystem cannot be broken even if the adversary has unlimited quantum computing power together with the best possible algorithms. Therefore, in theory, QKD generates highly secure cryptographic keys, immune to any algorithmic cryptanalysis, even those that may arise in quantum computing and which do threaten classical key-exchange systems [38].
Figure 2: Main scheme to deliver end-to-end quantum-secured cryptography using QKD. The symmetric quantum-distributed key is delivered to both ends of a cryptography system. For this, either a point-to-point QKD link exists, or an approach such as having trusted nodes that allow hop-by-hop key forwarding is used.
Since optical quantum channels suffer from a series of limitations regarding throughput and range, a network is needed to provide any significant end-to-end service. The quantum information is carried embedded in quanta of energy which, by their nature, are extremely sensitive to interactions. A widely used approach to achieve end-to-end key distribution is using “trusted nodes”, intermediate nodal sites where forwarding operations can be carried out without compromising the security. In each node, a component named “forwarding module” (FM) performs a key transport protocol, consuming the generated key and using another ITS primitive—e.g., Vernam encryption [39]—for enciphering and deciphering the final distributed key in each hop.
This enables end-to-end quantum-enhanced services both between remote subscribers and optical network itself—e.g., encryptors [40] for protecting the subscribing traffic—. In practice, cryptosystems based on QKD are not ITS. Any implementation requires assumptions which may not exactly reproduce a theoretical model. All elements of the network as a whole must be as secure as possible.
@@
Software-defined networking enables network programmability for enhancing the service provided [6, 42]. Traditionally, network equipment was design to save costs by limiting its configurability and focusing on performing specific tasks. Modern approaches promote the supply of more open and general-purpose equipment, often making use of software engineering techniques, such as exposing its functionality through software interfaces or domain-specific programming languages. This allows decoupling of functional planes, both the control and forwarding planes as well as the application plane; component disaggregation; and other software-enabled techniques such as virtualisation. In most cases, the SDN design assumes a logical centralisation of the control functions into a specific component called controller, which manages the underlying general-purpose infrastructure. The result is a network that can be programmed to meet the requirements of each application, either in concrete sense—e.g., each network usage per subscriber—or abstract—high-performance usage or low-energy impact usage—.
@@
Analogously, quantum networks can adopt this approach. In fact, this work discusses and applies the concept of the “quantum forwarding plane” (QFP): functionalities that enables the end-to-end quantum-enhanced provision. This includes the FMs that enable key forwarding in trusted nodes, as well as quantum information systems—e.g., QKD systems—and other infrastructure elements. All the elements in the QFP may be controlled by an unique control plane. This will be fundamental for the integration of quantum technologies. A proper and joint management of classical and quantum resources can address low-level issues, such as sharing the underlying infrastructure, as well as other issues already mentioned, such as security, legal requirements, and quality assurance.
@@
These and other issues are the subject of standardization, which is a fundamental task for QKD networks. The work described in this manuscript benefits from the standardisation activities on SDN in quantum networking [22, 23] driven by the ETSI GS QKD industry specification group.
@@
In this regard, the ETSI GS QKD 015 specification [43] defines a software-defined node in a QKD networks as a set of components, including two forwarding elements: the QKD modules, which handle quantum signals and protocols, and the set of local key management systems (LKMS), that securely manages and forwards the key. Fig. 3 depicts this architecture. This specification also determines an information model that enables controlling every nodal component in a software-defined way. The subscribers may request end-to-end QoS-driven QKD services by ETSI GS QKD 004 [44], passing the use-specific requirements to be fulfilled by the network. Other specifications, such as ETSI GS QKD 018 [45], enable inter-domain orchestration or key management inter-operation. The architecture of the network elements is also being standardized to include the definitions of the QFP and the FM used above [46].
@@ III-E Optical quantum networks
Quantum signals are of very low power and sensitive to interactions [17]. When transmitted in a shared optical fibre, the Raman effect and four-wave mixing of adjacent channels can cause disturbances in the quantum channels [51, 52]. In consequence, the management of the quantum and classical channel allocation is fundamental [53]. Active components, such as signal amplifiers, are destructive for the quantum nature of signals, regardless of the re-amplification, re-shaping or re-timing technique that they implement. The possibility of using other signal reconditions of passive nature, such as polarisation or dispersion compensation, may depend on the nature of the quantum communication.
In optical networking, multiplexers and other elements are not perfect and its effects, e.g., crosstalk between channels, signal reflections, and backscattering might influence the guardband communications [17, 54, 55]. As a result, coexistence must be implemented with optical equipment with good inter-channel isolation and limited launch power of the classical signals. Networks adopting the ROADM approach often use equipment that integrates amplification by design, such as wavelength selective switches (WSS). The quantum channels have to be dropped, filtered and added outside of the WSS to comply with the requirements. These restrictions have a direct impact on both the design and technology selection for the network.
When adopting such solutions, the requirements of the classical signals need to be managed for protecting the quantum channel. The power control negatively influence the propagation of classical signals. In order not to disturb a quantum system, it might be necessary to reduce the launch power of a classical channel so that the received power falls below the sensitivity of the classical receiver. The requirements could be met with better equipment or additional amplification on the receiver side once the channel has been dropped.
For many of these problems, as noted above, interoperability between the control and management functionalities of all elements becomes critical. If the operation and management system for the quantum resources can interoperate with that of the optical infrastructure, information can be consolidated and decisions can be made to increase performance. However, it is fundamental that industry-wide standards are developed to facilitate such coexistence and integration.

### v2

@@ I Introduction
The principles of quantum mechanics imply that the measurement of a quantum signal alters it. Passive networks, without active elements such as amplifiers, must be used for its transmission. The propagation of signals at the quantum level in physical media—typically the optical fibre—entails absorption probability that dramatically limits their range. If the medium is shared, other quantum or classical signals may limit the range even further. In the case of optical fibres, deployment is typically the most expensive part and co-propagation is essential. Until quantum repeaters are developed [3], it is necessary for QKD networks to handle these limitations.
Equally relevant to the success of quantum communications is the provision of quantum-enhanced services. This involves the joint operation and management of quantum and classical resources, which requires low-level enabling techniques allowing sharing the underlying physical infrastructure. To achieve widespread deployment, interoperability of deployed systems, standardisation, certification of quantum networks and devices, compliance with regulations, and quality assurance, are also a must. MadQCI [4] is designed to be heterogeneous at the quantum as well as at the classical network level, to comprehensively address these problems. It is also large and complex enough to significantly advance the state of the art. To date, with three years of continuous operation, MadQCI is the most complex, largest, and longest running testbed for quantum-communication technologies in Europe. To realise a diverse network and perform a comprehensive investigation, multiple technologies from various vendors were installed. The 26 QKD modules came from three different manufacturers, some using continuous-variable QKD (CV-QKD) and others discrete-variable QKD (DV-QKD). Table I summarises the main features of the network.
@@
Domain Nodes QKD links
Total link
length
QKD
modules
Encrypt.
modules
@@
The complexity of the network was tackled using several administrative domains. The infrastructure was hosted in production premises of two network operators: Telefónica Innovación Digital (TID), the research and development division of the incumbent telecommunications operator in Spain, and the research and education network operator in the region of Madrid, REDIMadrid (RM). In addition, a third (logical) domain was used to handle the testing of a switched QKD network. This specific sub-network is referred to as the “co-located domain” and contains all the devices that support switching [5, 6]. The network configuration resembles production networks—including the inter-domain border link—re-using already installed resources and adding new ones as required.
Regardless of the domain, the operation and management of the overall quantum network was based on a software-defined approach [4] to quantum networking. In the software-defined networking (SDN) paradigm, forwarding-plane resources are programmable and expose their capabilities through common and open interfaces [7]. In this way, the operation and management system not only configures and commands the resources, but also achieves the overall programming of the network for application-specific behaviour and performance. This design is very flexible and allows an easier integration of quantum resources in existing classical networks. In MadQCI, this paradigm enables the transparent operation and management of both the quantum and classical equipment, required for quantum-secure communications and services.
As a result, cryptographic keys could be distributed between any pair of nodes in the network, independently of the QKD manufacturer and network operator. These capabilities were exposed to applications through standardised interfaces, which benefited from the programmability of the network as a whole. The outcome allowed the execution of a wide variety of use cases [4, 8, 9] based on quantum-distributed keys. Specific hardware was installed to provide encryption at OSI layerss 1 and 2. Software applications allow secure communication on higher OSI layers, such as an IPsec implementation [10] and one based on TLS [11].
The paper is organised as follows. In the next sections, the related work and an overview about quantum and optical communications is outlined and the coexistence and integration approach of classical and quantum optical networks is discussed. Then, a detailed description of the solutions implemented in MadQCI and its main performance are presented to serve as a blueprint for large-scale quantum network deployments. The paper finishes with conclusions and an outlook.
@@ II Related work
Quantum cryptography has been the first major technological realisation of quantum communication and QKD grew around the idea of using optical communications and networks since its earliest days. Examples include the Boston testbed [12], which already involved optical networking techniques such as switching or dense wavelength division multiplexing (DWDM). Other well-known examples are the Vienna [13] and Tokyo [14] quantum networks, which used dark optical fibres deployed in the field. The first iterations of MadQCI also date back to that time [15], along with others such as the Hefei testbed [16].
Most of these practical exercises focused on showing the capabilities of quantum-communication developments, rather than on how to install these into existing infrastructures [17, 18, 19, 20, 21]. In contrast, the efforts in MadQCI have been aligned with this last goal since its early days. Novel optical networking approaches to quantum communication were proposed [22] and it was shown how a pioneering commitment to a software-defined approach for quantum networking enables a fast and flexible convergence of quantum and classical communication in production environments [23, 24]. Furthermore, the MadQCI team is a leading actor in the standardisation of QKD communication at ETSI ISG QKD [25], which promotes a dialogue between industrial actors such as manufacturers and network operators.
In parallel, numerous technological improvements over the last decade have brought QKD to a commercial status. Deployments such as the 2,000 km Beijing-Shanghai QKD link and further extensions of the China QKD Network [26, 27] as well as the 4,600 km long QKD space segment [28] have demonstrated the feasibility of quantum secured backbone networks. Other examples demonstrate the penetration of these technologies in specific niches, such as the ability of the Hefei quantum network to efficiently encrypt the institutional telephone traffic [29] or the efforts in Bristol to securely transport subscriber databases in 5G [30].
Since 2019, a broad effort is underway to establish the European Quantum Communications Infrastructure (EuroQCI), a pan-European quantum network based on the results of several projects funded by Quantum Flagship [31] and Digital Europe [32] research programmes, among others. Examples of these projects are OpenQKD [8], CIVIQ [33], and the current QSNP [34], which strongly contribute to results discussed in this paper. Currently, there are several efforts to deploy quantum communication infrastructures (QCI) in Europe. A list of these can be found in [35].
@@ III Quantum and optical communications overview
This section introduces some topics in telecommunications and quantum technologies, particularly those transmitting in optical fibres that are needed in the subsequent discussion.
@@ III-A Production networks
When considering the realization of quantum technologies in optical networks, it is necessary to build on a suitable model, which may differ depending on the nature of the requirements [38].
The current networking landscape relies on several heterogeneous communication systems with different design, technology, and service fundamentals. Besides, each of them has a different degree of interoperability, often limited by the commercial strategy of the respective vendors. Examples are the more traditional Ethernet-based local area networks; the passive optical networks (PONs) for home access; or the optical networks for aggregation and backhaul, as depicted in Fig. 1. All operate over optical fibres and can carry the same type of traffic. However, they have different requirements for the deployment of new systems. In addition, network operators and service providers must comply with legal and regulatory requirements—e.g. critical infrastructure protection, lawful interception, and secrecy of communications—that prevent the deployment of novel devices without certification and industry-wide standardisation. The deployment of quantum communication adds additional requirements to these already overloaded designs, all of which have to be integrated to deliver cutting-edge services.
@@
Different industrial and commercial approaches should also be taken into account when modelling the optical networks. Network operators focus their activities on the transport of information between their subscribers. Thus, their network design must deal with the complexity of equipment spread across customer sites and central offices, while assuring the aggregation and transport of multi-service traffic. Whereas, in the case of digital service providers, the complexity lies in their multi-tenant, distributed data centres, which are interconnected “over-the-top” of the network operator, abstracting the complexity of many parts of the network. The notion of quantum communications and the service they provide can be fundamentally different in these scenarios, and involves aspects ranging from regulatory issues to how or whom to bill for these services.
A description of a network can be obtained by outlining itsfunctionalities, usually classified in functional planes—i.e., the forwarding, control, management, application, etc. planes—. The previously explained context fixes how, when, and where those functionalities are executed. In this respect, a production network cannot be modeled as a naive graph of connected nodes. It is common for forwarding-plane functionalities, i.e., those functions that enable the transport of information over the different transmission media, to reside at all nodes of a network. However, the majority of the control-plane functionalities may reside in a single centralised component. In this case, the unavailability of a single node may compromise the connectivity of the entire graph. Furthermore, other elements, such as the facilities that physically host the forwarding devices and provide security and privacy to the system, are also functional to the network. Therefore, a more appropriate model should include which network functionalities are performed on which element, with a focus on the specific end use or application.
@@ III-B Optical networks
Finally, several versions or generations coexist in industry to achieve a higher or lower reconfigurability of lightpaths, as also depicted in Fig. 1. A common approach is called “FOADM”, which stands for fixed OADM. In these designs, elements such as fixed filters and multiplexers prevail, i.e., without the possibility of changing their frequency selectivity. Re-tunable transceivers may be used, depending on the connectivity requirements in the assembly. The “ROADM” approach uses precisely reconfigurable OADMs or analogous elements. This means that the optical mesh can be reconfigured by means of frequency selective elements to establish new lightpaths on demand. The degree of meshing may not be all-with-all, and hybrid approaches are possible—e.g., FOADM in the access network and ROADM in the backbone, or vice versa—.
@@ III-C Quantum key distribution and cryptographic applications
QKD provides synchronised sources of symmetric cryptographic keys between two remote systems [1, 2]. This technique relies on quantum physics to deliver information-theoretically secure (ITS) keys to a final cryptographic application [39], as Fig. 2 shows. Indeed, an ITS cryptosystem cannot be broken even if the adversary has unlimited (quantum) computing power together with ultimately efficient algorithms. Therefore, in theory, QKD generates unpredictable and unbreakable cryptographic keys, immune to any cryptanalysis, even such that may arise by replacing the communication channel by an adversarial quantum computer, or, as a corollary, by any quantum computing algorithms, which do threaten traditional key-exchange systems [40].
Figure 2: Main scheme to deliver end-to-end quantum-secured cryptography using QKD. The symmetric quantum-distributed key is delivered to both ends of a cryptography system. For this, either a point-to-point QKD link exists, or an approach, based on trusted nodes, that allows hop-by-hop key forwarding is used.
Since optical quantum channels suffer from a series of limitations regarding throughput and range, a network is needed to provide any significant end-to-end service. The quantum information is carried embedded in quanta of energy which, by their nature, are extremely sensitive to interactions. A widely used approach to achieve end-to-end key distribution is using “trusted nodes”, intermediate trusted nodal sites where forwarding operations can be carried out without compromising the security as a consequence of the stated trustworthiness of the locations. In each node, a component named “forwarding module” (FM) performs a key transport protocol, consuming the generated key and using another ITS primitive—e.g., Vernam encryption [41]—for ciphering and deciphering the final distributed key in each hop. (Additionally ITS message authentication is employed, as discussed e.g. in  [13].)
This enables end-to-end quantum-enhanced services both between remote subscribers and the optical network itself—e.g., by utilizing encryptors [42] for protecting the subscriber traffic—. In practice, cryptosystems based on QKD are not ITS. Any implementation requires assumptions which may not exactly reproduce a theoretical model. In any case, all elements of the network as a whole must be as secure as possible.
@@
Software-defined networking enables network programmability for enhancing the service provided [7, 44]. Traditionally, network equipment was designed to save costs by limiting its configurability and focusing on performing specific tasks. Modern approaches promote the supply of more open and general-purpose equipment, often making use of software engineering techniques, such as exposing its functionality through software interfaces or domain-specific programming languages. This allows decoupling of functional planes, both the control and forwarding planes as well as the application plane; component disaggregation; and other software-enabled techniques such as virtualisation. In most cases, the SDN design assumes a logical centralisation of the control functions into a specific component called controller, which manages the underlying general-purpose infrastructure. The result is a network that can be programmed to meet the requirements of each application, either in concrete sense—e.g., each network usage per subscriber—or abstract—high-performance usage or low-energy impact usage—.
@@
Analogously, quantum networks can adopt this approach. In fact, this work discusses and applies the concept of the “quantum forwarding plane” (QFP): functionalities that enables the end-to-end quantum-enhanced provision. This includes FMs that enable key forwarding in trusted nodes, as well as quantum information systems—e.g., QKD systems—and other infrastructure elements. All the elements in the QFP may be controlled by an unique control plane. This will be fundamental for the integration of quantum technologies in telecommunications. A proper and joint management of classical and quantum resources can address low-level issues, such as sharing the underlying infrastructure, as well as other already mentioned issues, such as security, legal requirements, and quality assurance.
@@
These and other topics are the subject of standardization, which is a fundamental task for QKD networks. The work described in this manuscript benefits from the standardisation activities on SDN in quantum networking [23, 24] driven by the ETSI GS QKD industry specification group.
@@
In this regard, the ETSI GS QKD 015 specification [45] describes a software-defined node in QKD networks as a set of components, including the QKD modules, which handle quantum signals and protocols, and the set of local key management systems (LKMS), that securely manage the key and invoke the (integrated) FMs that forward it. Fig. 3 depicts this design. This specification also determines an information model that enables controlling every nodal component in a software-defined way. The subscribers may request end-to-end QoS-driven QKD services by ETSI GS QKD 004 [46], passing the use-specific requirements to be fulfilled by the network. Other specifications, such as ETSI GS QKD 018 [47], enable inter-domain orchestration or key management inter-operation. The architecture of the network elements is also being standardized to include the definitions of the QFP and FM used above [48, 49].
@@ III-E Optical quantum networks
Quantum signals are of very low power and sensitive to extremely weak interactions [18]. When transmitted in a shared optical fibre, the Raman effect and four-wave mixing of adjacent channels can cause disturbances in the quantum channels [54, 55]. As a consequence, the management of the quantum and classical channels allocation is fundamental [56]. Active components, such as signal amplifiers, are destructive for the quantum nature of signals, regardless of the re-amplification, re-shaping or re-timing technique that they implement. The possibility of using other signal inverse reshaping passive methods, such as polarisation or dispersion compensation, may depend on the nature of the quantum communication.
In optical networking, multiplexers and other elements are not perfect and their effects, e.g., crosstalk between channels, signal reflections, and backscattering might influence the guardband communications [18, 57, 58]. As a result, coexistence must be implemented with optical equipment with good inter-channel isolation and limited launch power of the classical signals. Networks adopting the ROADM approach often use equipment that integrates amplification by design, such as wavelength selective switches (WSS). The quantum channels have to be dropped, filtered and added outside of the WSS to comply with the requirements. These restrictions have a direct impact on both the design and technology selection for the network.
When adopting such solutions, the requirements on the classical signals need to be managed for enforcing protection of the quantum channel. The power control negatively influences the propagation of classical signals. In order not to disturb a quantum system, it might be necessary to reduce the launch power of a classical channel so that the received power falls below the sensitivity of the classical receiver. These requirements could be met with better equipment or additional amplification on the receiver side once the channel has been dropped.
For many of the problems, noted above, interoperability between the control and management functionalities of all elements becomes critical. If the operation and management system for the quantum resources can interoperate with that of the optical infrastructure, information can be consolidated and decisions can be made to increase performance. However, it is fundamental that industry-wide standards are developed to facilitate such coexistence and integration.